OpenStudy (anonymous):

Please help with quadratics: Solve x2 + 6x + 7 = 0 1) x = -1 and x = -5 2) 3±2√2 3) −3±2√2 4) −3±2√2/ 2

OpenStudy (anonymous):

\[\left\{x\to -3-\sqrt{2},x\to -3+\sqrt{2}\right\} \]

OpenStudy (anonymous):

Solve 3x2 + 4x = 2. (2 points) Question 3 options: 1) 2 +or- √10 all over 6 2) -2 +or- √10 all over 3 3) -2 +or- √10 all over 3 4) -4 +or- √10 all over 3

OpenStudy (anonymous):

Using Mathematica, the solution for 3x2 + 4x = 2 is:\[\left\{x\to \frac{1}{3} \left(-2-\sqrt{10}\right),x\to \frac{1}{3} \left(-2+\sqrt{10}\right)\right\} \]When the solutions are expanded the following is the case:\[\left\{x\to -\frac{2}{3}-\frac{\sqrt{10}}{3},x\to -\frac{2}{3}+\frac{\sqrt{10}}{3}\right\} \]2) and 3) of your option list are identical and are the same as the first solution of the solution list above.